Two resistors, R1 = 200 Ω and R2 = 800 Ω are connected in series across a d.c. supply, Vs = 20 V

Calculate voltage drop V2 across R2 .
1
Expert's answer
2015-01-12T07:22:56-0500
Solution:
Given:
R1=200 Ohm
R2=800 Ohm
Vs=20V
Lets find currents I=I1=I2=Vs/(R1+R2)
But I2=V2/R2 -> V2=I2*R2=Vs*R2/(R1+R2)
I2= 16V
Answer: voltage drop V2 across R2 is 16V